about y' = x-y, y(0) = 0,

(1) find solution

(2) use direction field draw approximation solution on xy plane and compare with the result (1)
Answer #1

solution 1 [Page-t] Griven ODE with initial Condition : y = x-y, 700)=0-) 7 ay [linear Equation of Now, solving equ- (2) firsNow, Page-2] using Initial condition, y (0) = 0 from Equ. 3), we have. yo) = (0-1) + C. eo 7 col=0 Therefore solution is , Thاااااااا slope - Field :- کیا م عنی ہے oاورSlope - Field - - ( y= x+y14: ! ! ! } } } } } } } } } } } } } 5 45 4 35 35:54 :51 : : 3 4 5